OpenStudy (goformit100):

Check whether the following statement is true or not. If x , y ∈ Z are such that x and y are odd, then xy is odd.

terenzreignz (terenzreignz):

Where are you getting these? LOL

OpenStudy (goformit100):

@Noemi95

terenzreignz (terenzreignz):

Suppose \[\huge x,y \in \mathbb{Z}\] such that both x and y are odd... then there exist integers...\(\large h,k \in \mathbb{Z}\) such that \[\Large x = 2h+1\\\Large y = 2k+1\] Having said that... \[\Large xy = (2h+1)(2k+1) =4hk+2(h+k)+1\] \[\Large xy=2(\color{blue}{2hk+h+k})+1\] Since h and k are integers, \[\Large \color{blue}{2hk+h+k}\in \mathbb{Z}\]And therefore xy is odd ∎